# Watching, Reasoning and Searching: A Video Deep Research Benchmark on Open Web for Agentic Video Reasoning

### v1

@@ Abstract
Chengwen Liu1*, Xiaomin Yu2*, Zhuoyue Chang1*, Zhe Huang10*, Shuo Zhang10*, Heng Lian10, Kunyi Wang3,10, Rui Xu4, Sen Hu5,10, Jianheng Hou6, Hao Peng1, Chengwei Qin2,9, Xiaobin Hu7, Hong Peng1†, Ronghao Chen5,10†, Huacan Wang8,10† 1LZU, 2HKUST(GZ), 3UBC, 4FDU, 5PKU, 6USC, 7NUS, 8UCAS, 9HKUST, 10QuantaAlpha *These authors contributed equally to this work. †Correspondence: pengh@lzu.edu.cn, chenronghao@alumni.pku.edu.cn, wanghuacan17@mails.ucas.ac.cn Github: https://github.com/QuantaAlpha/VideoDR-Benchmark
@@ 1 Introduction
Based on these gaps, we propose VideoDR, the first open-domain benchmark that systematically evaluates video deep research. As shown in Figure 2, we extend deep research to an open-domain factoid question answering setting conditioned on video: models must extract and compose visual anchors from multiple frames Yang et al. (2025b); Chen et al. (2025a); Zhang et al. (2025), use browser-based search to locate candidate evidence on the open web Wei et al. (2025); Wu et al. (2025); Li et al. (2025b); Chen et al. (2025b); Zheng et al. (2025), and perform multi-hop reasoning in the joint evidence space of videos and webpages to output a unique and verifiable answer Liang et al. (2025). To support this setting, we explicitly incorporate an interactive web search process into the task definition and adopt annotation and strict quality control over diverse real-world scenarios to systematically remove samples that can be answered by the video alone or by webpages alone, making the combined capability of video understanding, web search, and evidence-based reasoning the core evaluation target Yao et al. (2022); Gao et al. (2023). Based on VideoDR, we evaluate mainstream models under both the Workflow and Agentic paradigms Liu et al. (2025), including closed-source models GPT-4o Hurst et al. (2024) and Gemini-3-pro-preview Team et al. (2023), as well as open-source models MiniCPM-V 4.5 Yu et al. (2025b), Qwen3-Omni-30B-A3B Xu et al. (2025), and InternVL3.5-14B Wang et al. (2025b), and conduct comprehensive performance and error analyses along three dimensions: difficulty, video duration, and semantic domain.
@@ 2 Related Work
Deep Research Benchmarks.
Existing deep research evaluations typically assess search, reasoning, and tool use as a unified process Wei et al. (2025); Wu et al. (2025); Li et al. (2025b); Zheng et al. (2025): one line of work tests multi-step query planning and information localization in live web environments Wei et al. (2025); Wu et al. (2025), while another studies more controllable settings to improve reliability and stability Xue et al. (2025); Chen et al. (2025b); Jin et al. (2025); Li et al. (2025a). Overall, however, most benchmarks still start from textual queries Wei et al. (2025); Wu et al. (2025), and visual content is often downplayed as static auxiliary information rather than first-class evidence in the retrieval and verification loop Jiang et al. (2024); Liang et al. (2025); Liu et al. (2025).
Video Reasoning Benchmarks.
Existing video reasoning evaluations are conducted under a closed-evidence setting, where questions are designed to be answerable using only the video itself, and they primarily stress long-video temporal understanding and long-context reasoning Fu et al. (2025); Li et al. (2024a); Wu et al. (2024); Zhou et al. (2025); Wang et al. (2025a); Fang et al. (2024); Yang et al. (2025a); Li et al. (2024b); Nagrani et al. (2024); Yu et al. (2025a). Recent agentic video efforts also explore multi-round interaction and tool calling within the video-understanding loop Zhang et al. (2025); Yang et al. (2025b); Chen et al. (2025a). In summary, systematic evaluations of multi-step evidence gathering and reasoning integration on the open web anchored to video cues remain relatively scarce.
@@ 3.1 Task Definition
We propose Video Deep Research (VideoDR): an open-domain factoid question answering benchmark conditioned on a given video, designed to evaluate a model’s ability to perform complex reasoning anchored in the video while leveraging the open web. Given a video and a natural language question , the model can interactively call a browser search tool , iteratively searching between video cues and web-page evidence, and finally output a factual answer . Each VideoDR sample is constructed such that the model must exploit multi-frame cues from the video to locate candidate evidence on the open web, and then perform multi-hop reasoning over the joint evidence space of the video and the open web in order to obtain a unique answer. Formally, the VideoDR task can be expressed as:
@@ 3.3.2 Human Testing
To verify the correctness of annotated questions and characterize their difficulty, we adopt a human evaluation protocol with five independent participants solving the tasks in a blind manner. For each sample , five subjects independently produce answers under the condition that they can only access the annotated video and the textual question, without being given any reference answers, forming the answer set . Subjects are required to complete the task in a browser environment and must, for each question, submit a final answer obtained via video browsing and autonomous web search. This process serves two purposes: ❶ to validate the solvability of the questions and the correctness of the annotations under conditions close to real-world usage; and ❷ to provide empirical evidence for subsequent difficulty stratification and estimation of the human upper bound.
@@
Under the above setting, we estimate the human upper bound by the average accuracy of the 5 subjects over all samples:
Figure 3(b) shows the human test results. Across all 100 samples, the mean sample-level success rate is with a standard deviation of . When results are stratified by difficulty, the mean is for Low, for Mid, and for High. Overall, human participants can reliably solve the majority of samples, and we do not observe a systematic pattern of universal failure, which indirectly supports the consistency of the annotated answers. In addition, during dataset construction, we manually re-examine samples with clear disagreement or samples where multiple mutually inconsistent yet seemingly reasonable answers are provided; if ambiguity remains after multiple rounds of verification, the sample is discarded. The final retained dataset ensures that each question corresponds to a unique and verifiable gold answer, given the available evidence, thereby reducing evaluation noise.
@@ 3.4 Data Statistics
Finally, we construct 100 VideoDR samples. To characterize the benchmark’s structural properties, and to provide a basis for stratified analyses in subsequent experiments, we summarize the dataset from three perspectives: ❶ video category distribution, ❷ question length distribution, and ❸ video duration distribution.
@@
As shown in Figure 4(a), VideoDR spans six semantic domains: Daily Life, Economics, Technology, Culture, History, and Geography. The distribution is relatively balanced: Daily Life accounts for 33%, Economics 16%, Technology and Culture 15% each, History 11%, and Geography 10%. This distribution ensures broad coverage of diverse content types in open-domain settings.
@@
As shown in Figure 4(c), on the video side, durations exhibit a clear long-tailed pattern: most videos cluster in a short-duration range, while a small number of videos longer than 10 minutes form the tail. This duration structure covers both rapid cue capture in short-video scenarios and cross-segment association in long-video scenarios, enabling a more comprehensive test of models’ cross-frame understanding and evidence localization across different time scales.
@@ 4.1 Setting
Baselines. We compare two common paradigms: Workflow and Agentic. Workflow adopts a two-stage design: a multimodal model first extracts question-relevant cross-frame visual cues from the video and produces a structured intermediate text, which is then provided together with the question as input for subsequent reasoning. Without further access to the original video, the system uses the search tool to retrieve candidate evidence from the open web, and uses the think tool to reflect on and filter evidence and to generate the next-round query, finally aggregating multi-round search results to produce an answer. In contrast, Agentic adopts a stronger end-to-end setting: it feeds the raw video and the question directly into a single multimodal agent, which performs video understanding, query generation, web retrieval, and evidence integration within the same execution loop. The agent likewise uses only the think tool and the search tool, and autonomously decides when to initiate search based on video cues, when to reflect, and how to organize multi-round evidence into the final conclusion.
MLLMs.
We select mainstream multimodal models spanning both closed-source and open-source families as the core research agents. Closed-source models include Gemini-3-pro-preview Team et al. (2023), GPT-4o Hurst et al. (2024), GPT-5Hurst et al. (2024); open-source models include MiniCPM-V 4.5 Yu et al. (2025b), InternVL3.5-14B Wang et al. (2025b), and Qwen3-Omni-30B-A3B Xu et al. (2025) . All models are tested under both Workflow and Agentic paradigms to analyze their capability boundaries under different system organizations.
LLM as Judge.
VideoDR is an open-domain factual question answering task. To avoid false mismatches, we adopt an LLM-as-judge protocol Zheng et al. (2023) using DeepSeek-V3-0324 Liu et al. (2024) to assess semantic equivalence between the model prediction and the reference answer. The judge outputs a binary correctness label, which we use to compute overall Accuracy.

### v2

@@ Abstract
Chengwen Liu1*, Xiaomin Yu2*, Zhuoyue Chang1*, Zhe Huang10*, Shuo Zhang10*, Heng Lian10, Jisheng Dang1†, Rui Xu4, Sen Hu5,10, Jianheng Hou6, Chengwei Qin2,9, Xiaobin Hu7, Kunyi Wang3,10, Zhi Yang10, Hao Peng1, Hong Peng1†, Ronghao Chen5,10†, Huacan Wang8,10† 1LZU, 2HKUST(GZ), 3UBC, 4FDU, 5PKU, 6USC, 7NUS, 8UCAS, 9HKUST, 10QuantaAlpha, *These authors contributed equally to this work. Project Leader: Xiaomin Yu yuxm02@gmail.com †Correspondence: pengh@lzu.edu.cn, dangjisheng@lzu.edu.cn, chenronghao@alumni.pku.edu.cn, wanghuacan17@mails.ucas.ac.cn Github: https://github.com/QuantaAlpha/VideoDR-Benchmark
@@ 1 Introduction
Based on these gaps, we propose VideoDR, the first open-domain benchmark that systematically evaluates video deep research. As shown in Figure 2, we extend deep research to an open-domain factoid question answering setting conditioned on video: models must extract and compose visual anchors from multiple frames Yang et al. (2025b); Chen et al. (2025a); Zhang et al. (2025), use browser-based search to locate candidate evidence on the open web Wei et al. (2025); Wu et al. (2025); Li et al. (2025b); Chen et al. (2025b); Zheng et al. (2025), and perform multi-hop reasoning in the joint evidence space of videos and webpages to output a unique and verifiable answer Liang et al. (2025). To support this setting, we explicitly incorporate an interactive web search process into the task definition and adopt annotation and strict quality control over diverse real-world scenarios to systematically remove samples that can be answered by the video alone or by webpages alone, making the combined capability of video understanding, web search, and evidence-based reasoning the core evaluation target Yao et al. (2022); Gao et al. (2023). Based on VideoDR, we evaluate mainstream models under both the Workflow and Agentic paradigms (Liu et al., 2025). The evaluated models include closed-source models Gemini-3.1-pro-preview (Google AI for Developers, 2026) and GPT-5.4 (OpenAI, 2026), as well as open-source models Qwen3.5-35B-A3B (Qwen Team, 2026), InternVL3.5-14B (Wang et al., 2025b), Gemma-4-31B-it (Google DeepMind, 2026), and GLM-4.6V-Flash (Z.ai, 2025). We conduct comprehensive performance and error analyses along three dimensions: difficulty, video duration, and semantic domain.
@@ 2 Related Work
Deep Research Benchmarks. Existing deep research evaluations typically assess search, reasoning, and tool use as a unified process Wei et al. (2025); Wu et al. (2025); Li et al. (2025b); Zheng et al. (2025): one line of work tests multi-step query planning and information localization in live web environments Wei et al. (2025); Wu et al. (2025), while another studies more controllable settings to improve reliability and stability Xue et al. (2025); Chen et al. (2025b); Jin et al. (2025); Li et al. (2025a). Overall, however, most benchmarks still start from textual queries Wei et al. (2025); Wu et al. (2025), and visual content is often downplayed as static auxiliary information rather than first-class evidence in the retrieval and verification loop Jiang et al. (2024); Liang et al. (2025); Liu et al. (2025).
Video Reasoning Benchmarks. Existing video reasoning evaluations are conducted under a closed-evidence setting, where questions are designed to be answerable using only the video itself, and they primarily stress long-video temporal understanding and long-context reasoning Fu et al. (2025); Li et al. (2024a); Wu et al. (2024); Zhou et al. (2025); Wang et al. (2025a); Fang et al. (2024); Yang et al. (2025a); Li et al. (2024b); Nagrani et al. (2024); Yu et al. (2025). Recent agentic video efforts also explore multi-round interaction and tool calling within the video-understanding loop Zhang et al. (2025); Yang et al. (2025b); Chen et al. (2025a). In summary, systematic evaluations of multi-step evidence gathering and reasoning integration on the open web anchored to video cues remain relatively scarce.
@@ 3.1 Task Definition
We propose Video Deep Research (VideoDR): an open-domain factoid question answering benchmark conditioned on a given video, designed to evaluate a model’s ability to perform complex reasoning anchored in the video while leveraging the open web. Given a video and a natural language question , the model can interactively call a browser search tool , iteratively searching between video cues and web-page evidence, and finally output a factual answer . During the subsequent research process, the model is not allowed to repeatedly re-watch the video. Each VideoDR sample is constructed such that the model must exploit multi-frame cues from the video to locate candidate evidence on the open web, and then perform multi-hop reasoning over the joint evidence space of the video and the open web in order to obtain a unique answer. Formally, the VideoDR task can be expressed as:
@@ 3.3.2 Human Testing
To verify the correctness of annotated questions and characterize their difficulty, we adopt a human evaluation protocol with five independent participants solving the tasks in a blind manner. For each sample , five subjects independently produce answers with access to the annotated video, the textual question, and autonomous web search, but without being given any reference answers, forming the answer set . Subjects are required to complete the task in a browser environment and must, for each question, submit a final answer obtained via video browsing and autonomous web search. This process serves two purposes: ❶ to validate the solvability of the questions and the correctness of the annotations under conditions close to real-world usage; and ❷ to provide empirical evidence for subsequent difficulty stratification and estimation of the human reference performance.
@@
Under the above setting, we estimate the human reference performance by the average accuracy of the 5 subjects over all samples:
Figure 3(b) shows the human test results. Across all 500 samples, the mean sample-level success rate is with a standard deviation of . When results are stratified by difficulty, the mean is for Low, for Mid, and for High. Overall, human participants can solve a substantial portion of the samples, and we do not observe a systematic pattern of universal failure, which indirectly supports the consistency of the annotated answers. In addition, during dataset construction, we manually re-examine samples with clear disagreement or samples where multiple mutually inconsistent yet seemingly reasonable answers are provided; if ambiguity remains after multiple rounds of verification, the sample is discarded. The final retained dataset ensures that each question corresponds to a unique and verifiable gold answer, given the available evidence, thereby reducing evaluation noise.
@@ 3.4 Data Statistics
Finally, we construct 500 VideoDR samples. To characterize the benchmark’s structural properties, and to provide a basis for stratified analyses in subsequent experiments, we summarize the dataset from three perspectives: ❶ video category distribution, ❷ question length distribution, and ❸ video duration distribution.
@@
As shown in Figure 4(a), VideoDR spans six semantic domains: Daily Life, Technology, Culture, History, Economics, and Geography. The distribution is relatively balanced: Daily Life accounting for 33.6%, Technology 22.6%, Culture 13.0%, History 11.8%, Economics 11.4%, and Geography 7.6%. This distribution ensures broad coverage of diverse content types in open-domain settings.
@@
As shown in Figure 4(c), the video durations exhibit a long-tailed distribution, covering both short clips and substantially longer videos. Based on our duration-based grouping used in later stratified experiments, the benchmark contains short, medium, and long videos in comparable proportions: 32.4% Short, 30.2% Medium, and 37.4% Long. This duration structure covers both rapid cue capture in short-video scenarios and cross-segment association in long-video scenarios, enabling a more comprehensive test of models’ cross-frame understanding and evidence localization across different time scales.
@@ 4.1 Setting
Baselines. We compare two common paradigms: Workflow and Agentic. Workflow adopts a two-stage design: a multimodal model first extracts question-relevant cross-frame visual cues from the video and produces a structured intermediate text, which is then provided together with the question as input for subsequent reasoning. Without further access to the original video, the system uses the search tool to retrieve candidate evidence from the open web, and performs explicit reflection and evidence filtering to generate the next-round query, finally aggregating multi-round search results to produce an answer. In contrast, Agentic adopts a stronger end-to-end setting: it feeds the raw video and the question directly into a single multimodal agent, which performs video understanding, query generation, web retrieval, and evidence integration within the same execution loop. The agent uses the search tool as its external interface to the open web, while autonomously deciding when to initiate search based on video cues, how to refine its queries through intermediate reasoning, and how to organize multi-round evidence into the final conclusion.
MLLMs. We select mainstream multimodal models spanning both closed-source and open-source families as the core research agents. Closed-source models include Gemini-3.1-pro-preview Google AI for Developers (2026) and GPT-5.4 OpenAI (2026); open-source models include Qwen3.5-35B-A3B Qwen Team (2026), InternVL3.5-14B Wang et al. (2025b), Gemma-4-31B-it Google DeepMind (2026), and GLM-4.6V-Flash Z.ai (2025). All models are tested under both Workflow and Agentic paradigms to analyze their capability boundaries under different system organizations.
LLM as Judge. VideoDR is an open-domain factual question answering task. To avoid false mismatches, we adopt an LLM-as-judge protocol Zheng et al. (2023) using DeepSeek-V3-0324 Liu et al. (2024) to assess semantic equivalence between the model prediction and the reference answer. The judge outputs a binary correctness label, which we use to compute overall Accuracy.